Circular No. 8573

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

CBAT@CFA.HARVARD.EDU (science)

URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/1983 N2 AND C/1984 Q1 (SOLWIND) Further to IAUC 8566, R. Kracht reports two additional Kreutz comets

found and measured by himself on SOLWIND images. Following are the observations of C/1983 N2, which appeared generally starlike and of mag 2–3, with a possible faint tail in the final images; B. G. Marsden provides the following parabolic orbital elements for C/1983 N2: T = 1983 July 7.89 TT, ω = 81^{o}.43, Ω = 359^{o}.55, i = 142^{o}.23 (equinox 2000.0), q = 0.0049 AU.

1983 UT  α_{2000}  δ_{2000}
July 7.425  6^{h}55.^{m}7  +20^{o}56′
7.432  6 55.9  +20 57
7.439  6 56.0  +20 58
7.446  6 56.0  +21 00
7.454  6 56.2  +21 02
7.491  6 56.7  +21 07
7.498  6 56.8  +21 08
7.506  6 56.8  +21 10
7.513  6 57.0  +21 12
7.520  6 57.1  +21 14
7.558  6 57.6  +21 20
7.565  6 57.7  +21 21
7.572  6 57.7  +21 23
7.579  6 57.8  +21 24
7.586  6 58.0  +21 26

Following are positions for C/1984 Q1, which had a faint tail and faded rapidly from mag ∼ 3 in the first few images. Marsden provides the following parabolic orbital elements for C/1984 Q1: T = 1984 Aug. 23.22 TT, ω = 81^{o}.04, Ω = 355^{o}.72, i = 144^{o}.14 (equinox 2000.0), q = 0.0049 AU.

1984 UT  α_{2000}  δ_{2000}
Aug. 22.923  9^{h}58.^{m}8  +10^{o}56′
22.930  9 59.0  +10 56
22.937  9 59.2  +10 57
22.944  9 59.4  +10 58
22.952  9 59.7  +10 59
22.989  10 01.0  +11 02
22.996  10 01.1  +11 02
23.011  10 01.7  +11 05
2005 July 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green